# A paradigm shift to assembly-like finite element model updating

### v2

@@ Abstract
In general, there is a mismatch between a finite element model of a structure and its real behaviour. In aeronautics, this mismatch must be small because finite element models are a fundamental part of the development of an aircraft and of increasing importance with the trend to more flexible wings in modern designs. Finite element model updating can be computationally expensive for complex structures and surrogate models can be employed to reduce the computational burden. A novel approach for finite element model updating, namely assembly-like, is proposed and validated using real experimental data. The assembly-like model updating framework implies that the model is updated as parts are assembled. Benchmarking against the classical global, or one-shot, approach demonstrates that the proposed method is more computationally efficient since it takes 20% fewer iterations to obtain convergence, also using fewer parameters for the model evaluations. Despite the increase in computational performance, the new approach retains the fidelity of the global approach.
@@ 1 Introduction
The FEMU task is important across all fields of engineering, finding applications such as damage detection [10], identification [11] and, from a different perspective, in topology optimisation [12]. FEMU is also crucial in aeronautics, where having a reliable FEM is pivotal in meeting certification requirements [13]. Further challenges have recently arisen for FEMU in aeronautics, in particular for aircraft wings, because the design paradigm is shifting towards lightweight materials and slenderer wings, in order to improve their aerodynamic efficiency [14]. Hence, developing accurate models is not only necessary for analysing aeroelastic effects, but also for design and controls. In particular, [15] develop a FEM of a scaled aircraft model to be updated with ground vibration testing (GVT) data from different loaded scenarios to estimate flutter onset speed. In [16] a FEMU technique is implemented for positioning a wing box composite material layers for passive aeroelastic suppression, while in [17] model updating is used for enhancing the handling qualities of an aircraft with flexible wings. FEMU has also been implemented for other applications, such as reverse engineering of a fighter aircraft internal structure [18]. Nevertheless, FEMU, in aeronautics, is not only carried out on wings, but also on the full aircraft, e.g. [19], and components, e.g. [20].
@@
Experimental data is required to update a FEM and in most cases, results from vibration data are used. The most prominent product of vibration data are modal parameters [21], which can be obtained by experiment, the so-called experimental modal analysis (EMA) [22], and during normal operation, known as operational modal analysis [23]. A particular type of EMA, used in aeronautics, is GVT and in this work, the two terms are used interchangeably as the work deals with aeronautical structures. In fact, modal parameters, in particular, natural frequencies ($\omega_{n}$) and mode shapes ($\bm{\phi}_{n}$) are a common metric for FEMU [24]. Notably, [25] used $\bm{\phi}_{n}$, via the modal assurance criterion (MAC) [26], from operational data for the FEMU of a wing section. Nevertheless, approaches based on the direct use of experimental data are still common, such as in [27] where a series of nonlinear case studies models has been updated using time domain responses.
@@
Having an accurate structural model of an aircraft structure is not fundamental only for the structural assessment per se, but also for the general life-cycle assessment of the aircraft. Recent advances in the digital twin paradigm have shown that this is vital for building reliable models [28]. These challenges are more challenging nowadays as wings become more flexible. For example, [29, 30] proposes a method for the numerical extraction of modal data from deformed wings (vertical displacement of the wing); however, this was not further explored experimentally. In [31] a component-based FEMU strategy is applied to a composite wing aircraft. Instead of updating the complete structure, the updating is carried out for an individual component, before assembly into the whole system; however, no comparison is given to one-shot approaches, neither in terms of performance nor precision. The latter is addressed, on the same system, in [32], showing favourable results, computationally speaking, for the progressive approach. However, the implementation suggested by Zhao in the aforementioned works, in particular their intermediate steps, is more similar to a hybrid testing approach [33, 34], as usually defined within Civil Engineering, where only part of a structure, or system (e.g. the port wing), is known, rather than to an assembly-like procedure. In fact, EMA data is available only for the wing and for the full assembly, rather than for all the components. Notably, a similar application is carried out for the structural health monitoring of a suspended bridge model in [35]. Moreover, a substructure-based finite element model updating approach is introduced in [36], wherein selected eigenmodes and derivatives from independent substructures are assembled into a reduced eigenequation, allowing global eigensolutions to be recovered with re-analysis restricted to the relevant substructures. Nevertheless, from these works it became apparent that the possibility of carrying out assembly-like model assembly, or updating, is appealing as one can consider smaller, in terms of computational loads (e.g. nodes and elements), models, rather than larger and more complex ones. However, a more dedicated approach to the paradigm shift from one-shot and partial model approaches to progressive assembly-like approaches is needed, since these works have not addressed the stepwise characterisation of all subassemblies.
@@
Given the aforementioned motivation, the novelty of this work is to propose and assess a framework for the assembly-like FEMU against the classical direct approach to show its goodness and efficiency. The framework will be tested on a flexible wing model, of which thorough experimental data, involving the full wing and its parts and sub-assemblies, is available in [37, 38]. The hypothesis is that there could be an improvement in model accuracy and computational performance compared to traditional one-shot approaches. The model updating procedure followed in this work is based on a surrogate-based iterative method based on rEGO, which aims at minimising the Modified Total Modal Assurance Criterion (MTMAC) [39] to match the discretised FEM modal response to experimental results. The rEGO is preferred over classical EGO as it provides superior computational efficiency, robustness, and precision [4]. Nevertheless, any other iterative optimiser and/or valid goal function could be implemented with the same concept.
The expected contribution of the work is to prove that the proposed, assembly-like FEMU, method is more computationally efficient when compared to the classical approach. Thus, the aims of this work are the following:
@@
    Obtain an updated discretised FEM of a complex structure.
2 The Refined Efficient Global Optimisation for Model Updating
rEGO [6, 4] was introduced to broaden the search capability of the Efficient Global Optimization (EGO) [40, 41]. The main aim was to establish rEGO as a global-local technique, in a hybrid sense [42], able to both navigate search spaces globally (avoiding local minima) and land, not only, in the area of the global minimum but as close as possible to it, via the implementation of refinement and selection techniques. The main novelties of EGO were the use of a Kriging surrogate model and the implementation of a new infill metric, the expected improvement (EI). The EI can be defined as a measure of how much could the known minimum improve if a given point was added to the data pool. The EGO workflow is similar to other surrogate-based techniques: (i) the design space is searched strategically, usually with Latin hypercube sampling [43], (ii) the absolute value of the EI is maximised to find the suitable infill point and (iii) the point is infilled. The process is iterated between (ii) and (iii) until convergence is reached. For EGO, usually, this happens when EI is less than 1% of the objective function minimum [40]. In order to improve this process, rEGO retains the same global structure but introduces two important principles: refinement and selection.
3 The Flexible Wing Model
The wing model chosen to validate the FEMU technique is a high aspect ratio (HAR) flexible wing model developed at Cranfield University for the Beam Reduction Dynamic Scaling (BeaRDS) project [45, 46, 47, 48]. The main aim of BeaRDS was to establish a workflow for the design, manufacture and testing of dynamically scaled HAR wings for use in Cranfield University 8$\times$6 ft (2.4$\times$1.8 m) wind tunnel. Specifically, the wing under scrutiny is the eXperimental BeaRDS-2 (XB-2) model (Figure 1(a)); a dynamically scaled model for an optimised (drag reduction) wing to be used in an A320-like aircraft. The scaled model is made up of four main components: the spar, the stiffening tube, the additional brass ballasts and the skin. The spar and stiffening tube serve as the wing torque box (Figure 1(b)), the brass ballasts were added for dynamic scaling purposes [49] and the skin was designed to minimise its effect on the wing stiffness [46]. Since dynamic scaling is not the focus of this work, the brass ballasts are removed for the current study. The spar is machined from two 6082-T6 Aluminium Alloy blocks and joined at mid-span with a weld and reinforced by four bolted L-section plates. Geometrically the spar can be divided into three main sections, according to its profile, as shown in Figure 1(a). The spar mass is 1.225 kg.
@@ 3 The Flexible Wing Model
The outer surface of the wing is defined by the skin (Figure 1(a)), made up of 47 different subsections, which are 3D printed in Digital ABS (white) and Agilus 30 (black), a rubber-like material. Despite the use of different materials for the adjacent strips, the skin is made only of three separate parts, thanks to the PolyJet technology, which allows for the 3D printing of subsequent layers with different materials [52]. This, and the use of a rubber-like material, gives enhanced flexibility to the wing skin, allowing for large tip displacements. The assembly of the torque box and the skin constitutes the full wing. Table 1 introduces the wing material and physical properties, including aerofoil, aspect ratio (AR), mean aerodynamic chord ($\bar{c}$), taper ratio ($\lambda$), Leading Edge sweep ($\Lambda_{\text{LE}}$) and quarter-chord sweep ($\Lambda_{c/4}$). Please note, the wing has a neutral twist and dihedral angles.
@@
Material E - Young Modulus [GPa] Poisson Ratio [-] $\bm{\rho}$ Density [kgm]-3]
@@
3.1 Preliminary Finite Element Models
@@
Table 3 shows properties of the sections of interest described above, including moments of inertia (Ixx and Iyy) in m4, areas (A) in m2 and torsion constants (J) in m4.
@@
Ixx [m4] 5$\times$10-8 6.50$\times$10-8 3.46$\times$10-8 1.29$\times$10-8 1.68$\times$10-10
Iyy [m4] 6.50$\times$10-7 2.30$\times$10-9 1.90$\times$10-9 1.24$\times$10-9 1.68$\times$10-10
A [m2] 14.60$\times$10-4 2.07$\times$10-4 1.69$\times$10-4 1.98$\times$10-4 1.49$\times$10-11
J [m4] 7$\times$10-7 4.5$\times$10-10 3.1$\times$10-10 3$\times$10-9 3.38$\times$10-10
@@ 3.2 Experimental Setup
3.3 Experimental Campaign and Preliminary Finite Element Models Results
@@
Notably, more modes are, generally, identified in the same frequency interval in the FEMs. Table 4 shows only the results for the modes identified between the first and last modes that can be found in both the experimental and FEMs data. At least three coincident modes are found for the spar (the first three bending modes) and torque box (the first bending and the first two coupled modes); however, this is not the case for the full wing. Only two coherent modes are identified and, thus, they are used for the FEMU process. Hence, three modes are considered for the FEMU of the spar and torque box, while only two for the FEMU of the full wing.
@@
4 Finite Element Model Updating of the Flexible Wing
@@ 4.4 The XB-2 wing
Model of XB-2 Torque box Skin Total
@@
These results highlight the computational advantage of the proposed assembly-like method and support the authors’ claims that a paradigm shift towards assembly-like methods needs to be considered in FEMU. On the other hand, this approach might not be suitable for non-assemblable structures, such as buildings (although, some new construction techniques could be considered to be assemblies and should be investigated). Nevertheless, other applicable examples suitable for assembly-like approaches could be small satellites, such as CubeSats, and other mechanical systems, such as aero-engines. However, complications might arise for systems that require altering the experimental setup to test their sub-components
@@ 5 Conclusions
    To the authors’ knowledge this is the first time that refined Efficient Global Optimisation is applied to a complex structure finite element model updating;
@@
Thus, the findings of this work support the paradigm shift of finite element model updating towards assembly-like approaches, which is the recommended approach for similar structures.

### v3

@@ Abstract
In general, there is a mismatch between a finite element model of a structure and its real behaviour. In aeronautics, this mismatch must be small because finite element models are a fundamental part of the development of an aircraft and of increasing importance with the trend to more flexible wings in modern designs. Finite element model updating can be computationally expensive for complex structures and surrogate models can be employed to reduce the computational burden. A novel approach for finite element model updating, namely assembly-like, is proposed and validated using real experimental data. The assembly-like model updating framework implies that the model is updated as parts are assembled. Benchmarking against the classical global, or one-shot, approach demonstrates that the proposed method is more computationally efficient since it takes 20% fewer iterations to obtain convergence, of which 95% comes from the subassembly models, which have fewer degrees of freedom and so are more computationally efficient. Despite requiring fewer model evaluations, the new approach retains the fidelity, within 1% of a joint natural frequencies and modal shapes index, of the global approach.
@@ 1 Introduction
The FEMU task is important across all fields of engineering, finding applications such as damage detection [10], identification [11], topology optimisation [12], and virtual testing [13]. FEMU is also crucial in aeronautics, where having a reliable FEM is pivotal in meeting certification requirements [14]. Further challenges have recently arisen for FEMU in aeronautics, in particular for aircraft wings, because the design paradigm is shifting towards lightweight materials and slenderer wings, in order to improve their aerodynamic efficiency [15]. Hence, developing accurate models is not only necessary for analysing aeroelastic effects, but also for design and controls. In particular, [16] develops a FEM of a scaled aircraft model to be updated with ground vibration testing (GVT) data from different loaded scenarios to estimate flutter onset speed. In [17] a FEMU technique is implemented for positioning a wing box composite material layers for passive aeroelastic suppression, while in [18] model updating is used for enhancing the handling qualities of an aircraft with flexible wings. FEMU has also been implemented for other applications, such as reverse engineering of a fighter aircraft internal structure [19]. Nevertheless, FEMU, in aeronautics, is not only carried out on wings, but also on the full aircraft, e.g. [20], and components, e.g. [21].
@@
Experimental data is required to update a FEM. In most cases, results from vibration data are used [22]. The most prominent product of vibration data are modal parameters [23], which can be obtained by experiment, the so-called experimental modal analysis (EMA) [24], and during normal operation, known as operational modal analysis [25]. A particular type of EMA, used in aeronautics, is GVT and in this work, the two terms are used interchangeably as the work deals with aeronautical structures. In fact, modal parameters, in particular, natural frequencies ($\omega_{n}$) and mode shapes ($\bm{\phi}_{n}$) are a common metric for FEMU [26]. Notably, [27] used $\bm{\phi}_{n}$, via the modal assurance criterion (MAC) [28], from operational data for the FEMU of a wing section.
@@
Having an accurate structural model of an aircraft structure is not fundamental only for the structural assessment per se, but also for the general life-cycle assessment of the aircraft. Recent advances in the digital twin paradigm have shown that this is vital for building reliable models [29]. These challenges become tougher nowadays as wings become more flexible. For example, [30, 31] propose a method for the numerical extraction of modal data from deformed wings (vertical displacement of the wing); however, this was not further explored experimentally. In [32] a component-based FEMU strategy is applied to a composite wing aircraft. Instead of updating the complete structure, the updating is carried out for an individual component, before assembly into the whole system; however, no comparison is given to one-shot approaches, neither in terms of performance nor precision. The latter is addressed, on the same system, in [33], showing favourable results, computationally speaking, for the progressive approach. However, the implementation suggested by Zhao in the aforementioned works, in particular their intermediate steps, is more similar to a hybrid testing approach [34, 35], as usually defined within Civil Engineering, where only part of a structure, or system (e.g. the port wing), is known, rather than to an assembly-like procedure. In fact, EMA data is available only for the wing and for the full assembly, rather than for all the components. Notably, a similar application is carried out for the structural health monitoring of a suspended bridge model in [36]. Moreover, a substructure-based finite element model updating approach is introduced in [37], wherein selected eigenmodes and derivatives from independent substructures are assembled into a reduced eigenequation, allowing global eigensolutions to be recovered with re-analysis restricted to the relevant substructures. Nevertheless, from these works it became apparent that the possibility of carrying out assembly-like model assembly, or updating, is appealing as one can consider smaller, in terms of computational loads (e.g. nodes and elements), models, rather than larger and more complex ones. However, a more dedicated approach to the paradigm shift from one-shot and partial model approaches to progressive assembly-like approaches is needed, since these works have not addressed the stepwise characterisation of all subassemblies.
@@
Given the aforementioned motivation, the novelty of this work is to propose and assess a framework for the assembly-like FEMU against the classical direct approach to show its goodness and efficiency.
The proposed framework differs from that proposed in [33], since mass and stiffness are addressed altogether for each sub-assembly. Additionally, the model updating strategy proposed follows a progressive approach (smaller to larger assembly), rather than the all-at-once updating approach, such that the sub-model and global model are updated at the same time. Furthermore, this is different from the substructure approach, because that in [37] only considers a substruture of the global model for FEMU, not all the sub-assemblies and final structure progressively, as proposed here.
The framework is tested on a flexible wing model, of which thorough experimental data, involving the full wing and its parts and sub-assemblies, is available [38, 39]. The hypothesis is that there could be an improvement in model accuracy and computational performance compared to traditional one-shot approaches. The model updating procedure followed in this work is based on a surrogate-based iterative method based on rEGO, which aims at minimising the Modified Total Modal Assurance Criterion (MTMAC) [40] to match the discretised FEM modal response to experimental results.Nevertheless, any other iterative optimiser and/or valid FEMU goal function could be implemented with the same concept.
The aim of the work is to prove that the proposed, assembly-like FEMU, novel method is more computationally efficient when compared to the classical approach. Thus, the objectives of this work are the following:
@@
    Obtain an updated discretised FEM of a complex model with the approach in (i).
2 Materials and Methods
2.1 The Flexible Wing Model
[…]
@@ 2.1 The Flexible Wing Model
The outer surface of the wing is defined by the skin (Figure 1(a)), made up of 47 different subsections, which are 3D printed in Digital ABS (white) and Agilus 30 (black), a rubber-like material. Despite the use of different materials for the adjacent strips, the skin is made only of three separate parts, thanks to the PolyJet technology, which allows for the 3D printing of subsequent layers with different materials [48]. This, and the use of a rubber-like material, gives enhanced flexibility to the wing skin, allowing for large tip displacements. The assembly of the torque box and the skin constitutes the full wing. Table 1 introduces the wing material and physical properties, including aerofoil, aspect ratio (AR), mean aerodynamic chord ($\bar{c}$), taper ratio ($\lambda$), Leading Edge sweep ($\Lambda_{\text{LE}}$) and quarter-chord sweep ($\Lambda_{c/4}$). Please note, the wing has neutral twist and dihedral angles.
@@
Material E - Young Modulus [GPa] Poisson Ratio [-] $\bm{\rho}$ - Density [kgm]-3]
@@
2.1.1 Preliminary Finite Element Models
@@
Table 3 shows properties of the beam-like sections of interest, including moments of inertia (Ixx and Iyy) in m4, areas (A) in m2 and torsion constants (J) in m4.
@@
[…]
@@ 2.1.2 Experimental Setup
2.1.3 Experimental Campaign and Preliminary Finite Element Models Results
@@
Notably, more modes are, generally, identified in the same frequency interval in the FEMs. Table 4 shows only the results for the modes identified between the first and last modes that can be found in both the experimental and FEMs data. At least three coincident modes are found for the spar (the first three bending modes) and torque box (the first bending and the first two coupled modes); however, this is not the case for the full wing. Only two coherent modes are identified and, thus, they are used for the FEMU process. Hence, three modes are considered for the FEMU of the spar and torque box, while only two for the FEMU of the full wing. It is worth noting that all coupled modes identified show couplig between bending and torsion.
@@
2.2 Finite Element Model Updating of the Flexible Wing
@@
For completeness, the selected optimisation routine (shown in Figure 3(a)) , rEGO, is discussed herein. Originally, rEGO [6, 4] was introduced to broaden the search capability of the Efficient Global Optimization (EGO) [58, 59]. The main aim was to establish rEGO as a global-local technique, in a hybrid sense [60], able to both navigate search spaces globally (avoiding local minima) and land, not only, in the area of the global minimum but as close as possible to it, via the implementation of refinement and selection techniques. The main novelties of EGO were the use of a Kriging surrogate model and the implementation of a new infill metric, the expected improvement (EI). The EI can be defined as a measure of how much could the known minimum improve if a given point was added to the data pool. The EGO workflow is similar to other surrogate-based techniques: (i) the design space is searched strategically, usually with Latin hypercube sampling [61], (ii) the absolute value of the EI is maximised to find the suitable infill point and (iii) the point is infilled. The process is iterated between (ii) and (iii) until convergence is reached. For EGO, usually, this happens when EI is less than 1% of the objective function minimum [58]. In order to improve this process, rEGO retains the same global structure but introduces two important principles: refinement and selection.
3 Results
@@ 3.3 The XB-2 wing
Furthermore, on the performance benchmarking, Figure 7 shows the MTMAC complement vs the iteration for the bottom-up and top-down cases. It is worth noting that the vertical dotted black lines and in-figure labels show the sub-assembly of reference of the bottom-up FEMU. Notably, more iterations are needed for the bottom-up approach. This might seem in opposition to what is stated about the number of model evaluations, which are lower for the bottom-up case. However, it should be recalled that at each iteration up to two model evaluations can be carried out when the $\epsilon_{1}$ is satisfied (see Figure 3). As expected, the MTMAC decreases as the number of iterations increases, and its final value is lower than the initial value for both FEMU strategies.
Figure 7: MTMAC complement value vs iteration.
@@
XB-2 Model Torque box Skin Total
@@
These results highlight the computational advantage of the proposed assembly-like method and support the authors’ claims that a paradigm shift towards assembly-like methods needs to be considered in FEMU. On the other hand, this approach might not be suitable for non-assemblable structures, such as buildings (although some new construction techniques could be considered to be assemblies and should be investigated). Nevertheless, other applicable examples suitable for assembly-like approaches could be small satellites, such as CubeSats, and other mechanical systems, such as aero-engines. However, complications might arise for systems that require altering the experimental setup to test their sub-components. Nevertheless, this approach is scalable to larger assemblies, as far as the structure and subsystems stay testable. A potential use case in aerospace could be that of smaller launch vehicles, for which it is difficult to obtain GVT data due to their size, but subassemblies are testable. Furthermore, this study has only included the modelling of the system linear dynamics, thus any nonlinearity in the system is not considered in the FEM or FEMU process. Finally, a change of the modelling strategy and optimisation goal function could potentially take these into account, but further validation would be needed as the superimposition principle is not directly applicable in the nonlinear regime.
@@ 4 Conclusions
    To the authors’ knowledge, this is the first time that refined Efficient Global Optimisation is applied to a complex structure finite element model updating;
@@
Thus, the findings of this work support the paradigm shift of finite element model updating towards assembly-like approaches, which is the recommended approach for similar structures, although its main limitation is the need for sub-assembly modal data.